OpenStudy (anonymous):

how to show the series diverges for the sum from k=2 to infinity k^(k-2)/3k

OpenStudy (anonymous):

\[\sum_{2}^{\infty} k ^{k-2}/3k\]

OpenStudy (amistre64):

one thought is to do the integral test

OpenStudy (anonymous):

but k on its power so i need to use ln?

OpenStudy (anonymous):

use the ratio test will be much easier

OpenStudy (amistre64):

\[\sum_{2}^{\infty} k ^{k-\cancel{2}^3}/3\cancel{k}\]

OpenStudy (anonymous):

i havent learned the ratio test yet =[

OpenStudy (amistre64):

1/3 pulls out and your left with \[1/3\sum_{2}^{\infty} k ^{k-3}\]

OpenStudy (amistre64):

this should be self evident that it blows up to infinity

OpenStudy (anonymous):

it should be K^(k-2) not K^(K-2^3) tho

OpenStudy (amistre64):

noone did k^k-2^3 thats a cancel and becomes a 3

OpenStudy (amistre64):

we have 1/k = k^-1 k^k-2-1=k^k-3

OpenStudy (anonymous):

should i use ln to (K-3) down first

OpenStudy (amistre64):

dunno, integral test was my first thought but i abandoned it for something a bit logical :)

OpenStudy (amistre64):

hitting post is like playing the lotto, will it freeze or post? lol

OpenStudy (anonymous):

loll what do you mean

OpenStudy (amistre64):

just trying to work the site. i go to hit post to respond and my system will freeze up sometimes

OpenStudy (anonymous):

i don't have the problem, its probably because this site updates every second

OpenStudy (amistre64):

but yeah, if you can show k^(k-3) does off into infinity then it diverges by definition

OpenStudy (amistre64):

goes off into ...